1) A price decreased from $60 to $45. Find the percent of decrease.
2) The price in dollars p minus a 15% discount is $3.50. find p

Answer by stanbon(75887)

1) A price decreased from $60 to $45. Find the percent of decrease.
Quantity decrease = $15
Percent decrease = 15/60 = 25%
----------------------------------------
2) The price in dollars p minus a 15% discount is $3.50. find p
p - 0.15p = 3.50
---
p(0.85) = 3.50
---
p = 3.50/(0.85)
---
p = $4.12
=============
Cheers,
Stan H.